OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 2 > -3 and 3x <= 12 Answer A number line with a closed circle on -1, shading to the left, and an open circle on 4, shading to the right. A number line with an open circle on -1, a closed circle on 4, and shading in between. A number line with an open circle on -1, shading to the left, and a closed circle on 4, shading to the right. A number line with a closed circle on -1, an open circle on 4, and shading in between.

OpenStudy (anonymous):

that helps... but which one is it?

OpenStudy (anonymous):

Sorry, old guy, don't know anything about circles and whatnot, does closed mean includes and open excluded?

jimthompson5910 (jim_thompson5910):

x - 2 > -3 and 3x <= 12 x > -3+2 and x <= 12/3 x > -1 and x <= 4 -1 < x and x <= 4 -1 < x <= 4 So this means that the graph on the number line looks like a line shaded segment from -1 to 4 but there's an open circle at -1 and a closed circle at 4